SOLVE 7x - 1 -2(x + 5) = - (x - 3)
Answer by stanbon(75887)

SOLVE 7x - 1 -2(x + 5) = - (x - 3)
-----
7x -1 -2x-10 = -x+3
---
5x-11 = -x+3
6x = 14
x = 7/3
=============
Cheers,
Stan H.
=============